OpenStudy (anonymous):

Emily wants to hang a painting in a gallery but the painting and frame must have an area of 31 square feet. The painting is 5 feet wide by 6 feet long. The frame must have an equal width of x on each side. Which quadratic equation represents the area of the painting and frame combined? 4x2 + 22x − 1 = 0 4x2 + 22x + 31 = 0 x2 + 11x − 1 = 0 x2 + 11x + 31 = 0

OpenStudy (anonymous):

x is the width of the frame. x has to be added twice to each dimension SOO the dimensions with the frame are (5 + 2x) and (6 + 2x) The area is (5 + 2x)(6 + 2x) and must be = 31 30 + 10x + 12x + 4x² = 31 4x² + 22x + 30 - 31 = 0 4x² + 22x - 1 = 0

OpenStudy (anonymous):

Thank you